OpenStudy (anonymous):

Suppose f(π/3) = 2 and f '(π/3) = −5, and let g(x) = f(x) sin x and h(x) = (cos x)/f(x). Find the following: g'(π/3) h'(π/3)

hartnn (hartnn):

g(x) = f(x) sin x whats g'(x) from this ?

OpenStudy (anonymous):

f'(x)sin(x)+f(x)sin'(x) I get it (5radical3)/2 +1 it keeps saying that it is wrong

hartnn (hartnn):

f'(x)sin(x)+f(x)sin'(x) whats (sin x)' = .... ?

OpenStudy (anonymous):

cos(x)

hartnn (hartnn):

f'(x)sin(x)+f(x)cos x = -5 sin(pi/3) +2 cos(pi/3) = -5 sqrt 3/ 2 +2 (1/2) =-5/2 sqrt 3 +1 you are actually correct......hmmm

hartnn (hartnn):

oh wait!

hartnn (hartnn):

you missed the -*minus* sign in the beginning :P